OpenStudy (anonymous):

What is the equation of the line that has a slope of 4 and a y-intercept of -7? If you answer correct i will become ur fan

OpenStudy (anonymous):

a.y = -7x + 4 b.y = 7x - 4 c.y = -4x + 7 d.y = 4x - 7

OpenStudy (anonymous):

use this formula y=mx+b m= slope, b= y intercept and plug in y=4x-7

OpenStudy (anonymous):

thank you